OpenStudy (anonymous):

If possible, solve each system of equations. Use any method. If there is not a unique solution to a system , state a reason. 3x-11y=9 -9x + 33y =18 Need the steps please.

OpenStudy (anonymous):

ok, which method you want to use? elimination or substitution?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

ok let's use elimination first 3x-11y=9 x3 -> 9x-33y=27 -9x + 33y =18 do addition, 9x-9x-33y+33y=27+18 0=45 both x and y are eliminated, so the system doesn't have solutions

OpenStudy (anonymous):

no unique solution as both variables are eliminated ie one equation is a product/multiple of the other not a true system

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

now we'll do substitution 3x-11y=9 3x= 9+11y x=3 + 11/3y -9x + 33y =18 -9(3 + 11/3 y) + 33y = 18 -18 -33y+33y = 18 we can't find y, so we also can't find x this prove that the system doesnt have solution

OpenStudy (anonymous):

Thanks.